Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

1-P method doesnt work.Please help

Expert replies
by dddanny2006 » Mon Nov 25, 2013 11:14 am
If the probabilities are respectively 0.09,0.15,.21 and 0.23 that a person purchasing a new automobile will choose the color green,white,red or blue,what is the probability that a given buyer will purchase a new automobile that comes in one of those colors.

Answer 0.68---------------Simple method

Method 1


Prob(of choosing atleast one of those cars)=1-prob(of not choosing atleast one of those cars)

Prob(of not choosing atleast one of those cars)= 0.91*0.85*0.79*0.77=0.48

There 1-P=0.52

Why am I wrong here?
Answer is 0.68


Thanks

Dan
Join the discussion
Source: — Problem Solving |

by [email protected] » Mon Nov 25, 2013 3:09 pm
Hi dddanny2006,

In this question, we're told that the probability of choosing a car in a certain color is:

.09 = green
.15 = white
.21 = red
.23 = blue
.32 = some other color (this is implied, since total probability always adds up to 1)

We're asked for the probability of picking a car in ANY of those 4 colors (meaning green OR white OR red OR blue).

The probability formula = (What we want) / (Total possibilities)

Since we want "green or white or red or blue", we can simply add those up: .09 + .15 + .21 + .23 = .68

.68/1 = .68

The probability of NOT getting one of those colors is .32/1 = .32
So, 1 = .32 = .68

You can get the same answer with either approach, as long as you're doing the calculation that the question asks for.

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by dddanny2006 » Mon Nov 25, 2013 3:21 pm
What wrong did I do in the question,or in other words what should have been the question for the answer I got via my method

Thanks
[email protected] wrote:Hi dddanny2006,

In this question, we're told that the probability of choosing a car in a certain color is:

.09 = green
.15 = white
.21 = red
.23 = blue
.32 = some other color (this is implied, since total probability always adds up to 1)

We're asked for the probability of picking a car in ANY of those 4 colors (meaning green OR white OR red OR blue).

The probability formula = (What we want) / (Total possibilities)

Since we want "green or white or red or blue", we can simply add those up: .09 + .15 + .21 + .23 = .68

.68/1 = .68

The probability of NOT getting one of those colors is .32/1 = .32
So, 1 = .32 = .68

You can get the same answer with either approach, as long as you're doing the calculation that the question asks for.

GMAT assassins aren't born, they're made,
Rich
Join the discussion